OpenStudy (anonymous):

The density of mercury is 13.6 g/mL. How many pounds does one quart of mercury weigh? (1 pound is 454 g, 4 quarts are 3.7854 liters.)

OpenStudy (anonymous):

13.6/454 = 0.03 lbs (3.7854/4) = 0.946 L = 1 qt 0.03 * 0.946 * 1000 = 28.37 lbs/qt

OpenStudy (anonymous):

Let me know if you need further explanation.

OpenStudy (anonymous):

no thats great :) thank you

OpenStudy (anonymous):

@McSpazzy are you good in earth space science??

OpenStudy (anonymous):

Not my best field...but not my worst =D

OpenStudy (anonymous):

until wat time ur goin to be online?? i want to take an exam n the subject is about the universe

OpenStudy (anonymous):

@doubledrive For a while.